Find proj,u. |v = 21 - 2] - 4k, u = 51 - 12k 04.쯤름쯤
Answer #1

V= 21-23-4k ; 4= 57 -124 projuu= UV v (012 = (2,-2,-4). 159-12) C. (21-2;- 46) (12,4 1-23 (-4) 2) ² = 215)-2002 -41412) (di-2